Solve for x:
x^2+5x=6(x+2)

Answer by tutorcecilia(2105)

x^2+5x=6(x+2)
x^2+5x=6x+12
x^2+5x-6x=6x-6x+12
x^2+(-x)-12=+12-12
x^2-x-12=0
(x-4)(x+3)
x=4 or x=-3.
Check by plugging the values for x back into the original equation.